# Evaluating many-body stabilizer Rényi entropy by sampling reduced Pauli strings: singularities, volume law, and nonlocal magic

### v2

@@ IV.2 Thermodynamic integration
$\begin{split}\frac{d\log Z(\lambda)}{d\lambda}=&-\beta\bigg{\langle}\frac{dE(\lambda)}{d\lambda}\bigg{\rangle}_{\lambda}\end{split}$ (22)
@@
$\log\frac{Z_{B}}{Z_{A}}=\int_{\lambda_{A}}^{\lambda_{B}}-\beta\bigg{\langle}{dE(\lambda)}\bigg{\rangle}_{\lambda}$ (23)
@@
If we restrict the intermidiate process of the integration to be adiabatic, repeated thermalization procedures can be saved by taking the final state from the last simulation as the initial state for the next simulation. This is a trick of simulated annealing [76] and also applies to the ReAn method. In this work, we use TI to calculate the derivatives of $2$-SRE by identifying $Z_{B}\equiv Q(\lambda_{1})$ and $Z_{A}\equiv Q(\lambda_{2})$ (similarly for $Z$ and $Z_{2}$). We use ReAn to compute the values of $\tilde{M}_{2}$ instead of TI because empirically TI requires more interpolations than ReAn, as the former one has to consider numerical integration errors.
According to Eq. (18), the first-order derivatives of the $2$-SRE with paramter $J$ and $\beta$ can be easily achieved as
$\frac{d\tilde{M}_{2}(J)}{dJ}=\langle E\rangle_{J,Q}-2\langle E\rangle_{J,Z}-\langle E\rangle_{J,Z_{2}}$ (24)
$\frac{d\tilde{M}_{2}(\beta)}{d\beta}=\langle E\rangle_{\beta,Q}-2\langle E\rangle_{\beta,Z}-\langle E\rangle_{\beta,Z_{2}}$ (25)
@@
$\frac{d^{2}\tilde{M}_{2}(J)}{dJ^{2}}=-\frac{1}{J^{2}}\bigg{[}C_{J,Q}-2C_{J,Z}-C_{J,Z_{2}}\bigg{]}$ (28)
$\frac{d^{2}\tilde{M}_{2}(\beta)}{d\beta^{2}}=-\frac{1}{\beta^{2}}\bigg{[}C_{\beta,Q}-2C_{\beta,Z}-C_{\beta,Z_{2}}\bigg{]}$ (29)
@@
$\begin{split}&\langle C\rangle_{Q,J}:=\langle n_{J,\text{tot}}(n_{J,\text{tot}}-1)\rangle_{J,Q}-\langle n_{J,\text{tot}}\rangle^{2}_{J,Q}\\ &\langle C\rangle_{Z,J}:=\langle n_{J,\text{tot}}(n_{J,\text{tot}}-1)\rangle_{J,Z}-\langle n_{J,\text{tot}}\rangle^{2}_{J,Z}\\ &\langle C\rangle_{Z_{2},J}:=\langle n_{J,\text{tot}}(n_{J,\text{tot}}-1)\rangle_{J,Z_{2}}-\langle n_{J,\text{tot}}\rangle^{2}_{J,Z_{2}}\\ \end{split}$ (30)
@@
$\begin{split}&\langle C\rangle_{Q,\beta}:=\langle n_{\text{tot}}(n_{\text{tot}}-1)\rangle_{\beta,Q}-\langle n_{\text{tot}}\rangle^{2}_{\beta,Q}\\ &\langle C\rangle_{Z,\beta}:=\langle n_{\text{tot}}(n_{\text{tot}}-1)\rangle_{\beta,Z}-\langle n_{\text{tot}}\rangle^{2}_{\beta,Z}\\ &\langle C\rangle_{Z_{2},\beta}:=\langle n_{\text{tot}}(n_{\text{tot}}-1)\rangle_{\beta,Z_{2}}-\langle n_{\text{tot}}\rangle^{2}_{\beta,Z_{2}}\\ \end{split}$ (31)
@@ V.2 Contributions from $Q$ and $Z$ for the singularities in the ground-state magic
The results above suggest that in general models, the interplay between the $Q$- and $Z$-parts, or the contributions from the characteristic function and free energy, may result in more intricate behavior of $\tilde{m}_{2}$. The the order of the phase transition plays a significant role here. From this perspective, it is more clear that magic quntified by $\tilde{m}_{2}$ does not necessarily achieve a maximum or minimum at the quantum critical point consequently. This argument is also supported by previous small-scale numerical results for the 2D $\mathbb{Z}_{2}$ gauge theory, obtained using tree tensor network sampling [11]. For the 1D and 2D TFI models studied here, the divergence of the $Q$-part in $[d^{2}\tilde{m}_{2}/dJ^{2}]$, which relates to $\Xi_{P}(\rho)$, suggests a more direct connection between magic and (conformal) criticality. However, this relationship is still subtle, as the $Q$-part may lack a singularity for certain models and general bases.
@@ V.3 Volume-law corrections
To gain deeper insight into the results and many-body magic, we discuss the volume-law corrections of $2$-SRE. For ground states that are product states, the strict volume law follows from the additivity property described in Sec. II.2. Meanwhile, any nonzero correction signifies the existence of magic residing in correlations, which cannot be removed via local non-Clifford operations. These corrections have similar role with the mutual magic $\mathcal{L}_{AB}(\rho):=\mathcal{M}(\rho_{AB})-\mathcal{M}(\rho_{A})-\mathcal{M}(\rho_{B})$ [9, 54, 15, 55, 56] in a bipartite system ${A}\cup{B}$, where $\mathcal{M}$ is some mixed-state magic montone. Additionally, the volume-law corrections are beyond the bipartition scenario.
To fit the nonlocal corrections, we write
@@
for ground states on a $d$-dimensional lattice with length $L$ with the data of $\tilde{m}_{2}$ in Fig. 2. The volume-law coefficients $a_{1}$ and $a_{2}$ for the TFI models are also shown in Fig.  2, which are close to their corresponding $2$-SRE density, confirming that volume law is exactly the leading term.
Fig. 6(a) presents the correction $b_{1}$ for the ground states of the 1D TFI model. Since both the paramagnetic phase and the ferromagnet phase are gapped, in the thermodynamic limit, no nonlocal term should be captured by $2$-SRE. This reasoning also extends to the 2D case shown in Fig. 6(b). Therefore, we attribute the nonzero values of $b_{d}$ ($d=1,2$) near $J_{c}$ to finite-size effects. More interestingly, these finite-size effects lead to the indication of discontinuities at the critical points. The difference between the two phases are also reflected by the discontinuities of the magical structures reflected by $b_{d}$. This phenomenon contrasts strongly with the full-state magic, from which extracting information of phases and criticalities proves challenging. Though the full-state magic is basis-dependent, the nonlocal correlations at the critical point could always require nonlocal non-Clifford operations. Therefore, we claim that the volume-law corrections of magic would provide a stronger diagnostic for criticalities compared to full-state magic. This is especially significant for $\alpha$-SRE, which, as a non-mixed-state measure, fails to provide a reliable or monotonic mutual magic for bipartite systems. By contrast, studying volume-law corrections provide a practical means of quantifying nonlocal magic and criticality with $\alpha$-SRE.
@@ V.4 $2$-SRE of 2D finite-temperature Gibbs states
At the infinite-temperature limit ($\beta=0$), $\tilde{m}_{2}$ is approximately zero within the error bar. For finite temperatures, however, $\tilde{m}_{2}$ exhibits complex behavior, offering minimal valuable information, as shown in Fig. 7(a). Additionaly, for the first-order derivative $d\tilde{m}_{2}/d\beta$, we observe a singular point $\beta^{*}\approx 0.687<\beta_{c}$ in the paramagnetic phase, as shown in Fig. 7(b).
@@
In this work, we introduce a novel and efficient QMC method for computing the $\alpha$-SRE and its derivatives in large-scale and high-dimensional quantum many-body systems. Our method is broadly applicable to any sign-problem-free Hamiltonian regardless of the dimension of the system. A key benefit of our approach is that it allows us to isolate the free energy contribution in $\alpha$-SRE and the derivatives, providing deeper understanding into magic and criticality in quantum many-body systems, which are inaccessible before. Another significance of our algorithm is that it shows not only certain magic states of a high-dimensional many-body Hamiltonian can be simulated efficiently on a classical computer, but their magic can be computed efficiently. This underscores the fact that magic is just a necessary but not sufficient resource for quantum advantage.
@@
Physically, we demonstrate our approach on 1D and 2D ground states as well as finite-temperature Gibbs states of the TFI model. We reveal that the behavior of the $\alpha$-SRE is governed by the interplay between the (logarithmic of) generalized partition function $Q$, related to the characteristic function, and the partition function $Z$, related to the free energy. This explains why full-state magic provides little insight into phases and criticality in general, as observed in previous studies. Meanwhile, the order of the phase transition plays a crucial role for the behaviors of $\alpha$-SRE and the singularities in its derivatives. This is quite different from quantum entanglement: for both the 1D and 2D TFI model, previous results show that the quantum entanglement reaches its maximum at the quantum critical points [81, 82].
@@
To further investigate the use of full-state magic, we consider the volume-law corrections and analyze the nonlocal magic inherent in system correlations. These corrections, observed in finite system sizes, show the evidence of discontinuity at quantum critical points, indicating distinct magical structures on the two sides of the critical point. We consider volume-law corrections to be more significant, and their study does not require a mixed magic monotone, unlike the mutual magic $\mathcal{L}_{AB}$ introduced in Sec. V.3. We will explore these corrections further for more systems in future work. Finally, we confirm that $\alpha$-SRE is not an effective measure for mixed-state magic in many-body systems by studying an example of finite-temperature phase transition.
@@
In closing, we discuss extending our algorithm to compute the magic of a reduced density matrix. The mutual magic $\mathcal{L}_{AB}$ quantifies those magic generates from nonlocal non-Clifford operations, and it would encode crucial information of many-body systems. Previous studies have shown that the mutual magic $\mathcal{L}_{AB}$ would exhibit stronger signatures than the full-state magic in detecting criticalities and is capable of characterizing certain states in topological quantum field theory [9, 54, 15, 11, 83, 55, 56]. While $\mathcal{L}_{AB}$ is ideally defined using a mixed-state magic monotone, using $\alpha$-SRE to define it, called the mutual stabilizer Rényi entropy (mSRE), has also proven valuable [11, 56, 83], much like quantum mutual information captures entanglement. Our algorithm can be easily extended to compute reduced density matrix magic by tracing out one subsystem in QMC simulations using the technique in [84, 85, 86, 87], so that the mSRE can be computed. This is also a promising way to study the finite-temperature phase transition and open quantum systems rather than considering the full-state $\alpha$-SRE, as mentioned at the beginning. We leave the exploration of the mSRE for future work, with the hope that tools from quantum information, particularly the complexity theory, will illuminate new aspects of quantum many-body physics. Our method will enhance the further fusion and intersection of quantum information and many-body physics.
@@ Acknowledgements
We thank P.S. Tarabunga, M. Dalmonte, E. Tirrito, H. Timsina, T. Byrnes, C. Radhakrishnan, K. Warmuz, Z. Liu, G.-Y. Zhu, and Z.-Y. Wang for helpful discussions. Z.W. acknowledges the China Postdoctoral Science Foundation under Grant No. 2024M752898. This work is supported by the start-up funding of Westlake University. The authors thank the high-performance computing center of Westlake University and the Beijing PARATERA Tech Co., Ltd. for providing HPC resources.

### v3

@@ IV.2 Thermodynamic integration
$\begin{split}\frac{d\log Z(\lambda)}{d\lambda}=&-\beta\bigg{\langle}\frac{dE(\lambda)}{d\lambda}\bigg{\rangle}_{\lambda}.\end{split}$ (22)
@@
$\log\frac{Z_{B}}{Z_{A}}=\int_{\lambda_{A}}^{\lambda_{B}}-\beta\bigg{\langle}{dE(\lambda)}\bigg{\rangle}_{\lambda}.$ (23)
@@
If we restrict the intermediate process of the integration to be adiabatic, repeated thermalization procedures can be saved by taking the final state from the last simulation as the initial state for the next simulation. This is a trick of simulated annealing [83] and also applies to the ReAn method. In this work, we use TI to calculate the derivatives of $2$-SRE by identifying $Z_{B}\equiv Q(\lambda_{1})$ and $Z_{A}\equiv Q(\lambda_{2})$ (similarly for $Z$ and $Z_{2}$). We use ReAn to compute the values of $\tilde{M}_{2}$ instead of TI because empirically TI requires more interpolations than ReAn, as the former one has to consider numerical integration errors.
According to Eq. (18), the first-order derivatives of the $2$-SRE with parameter $J$ and $\beta$ can be easily achieved as
$\frac{d\tilde{M}_{2}(J)}{dJ}=\langle E\rangle_{J,Q}-2\langle E\rangle_{J,Z}-\langle E\rangle_{J,Z_{2}},$ (24)
$\frac{d\tilde{M}_{2}(\beta)}{d\beta}=\langle E\rangle_{\beta,Q}-2\langle E\rangle_{\beta,Z}-\langle E\rangle_{\beta,Z_{2}},$ (25)
@@
$\frac{d^{2}\tilde{M}_{2}(J)}{dJ^{2}}=-\frac{1}{J^{2}}\bigg{[}C_{J,Q}-2C_{J,Z}-C_{J,Z_{2}}\bigg{]},$ (28)
$\frac{d^{2}\tilde{M}_{2}(\beta)}{d\beta^{2}}=-\frac{1}{\beta^{2}}\bigg{[}C_{\beta,Q}-2C_{\beta,Z}-C_{\beta,Z_{2}}\bigg{]},$ (29)
@@
$\begin{split}&\langle C\rangle_{Q,J}:=\langle n_{J,\text{tot}}(n_{J,\text{tot}}-1)\rangle_{J,Q}-\langle n_{J,\text{tot}}\rangle^{2}_{J,Q},\\ &\langle C\rangle_{Z,J}:=\langle n_{J,\text{tot}}(n_{J,\text{tot}}-1)\rangle_{J,Z}-\langle n_{J,\text{tot}}\rangle^{2}_{J,Z},\\ &\langle C\rangle_{Z_{2},J}:=\langle n_{J,\text{tot}}(n_{J,\text{tot}}-1)\rangle_{J,Z_{2}}-\langle n_{J,\text{tot}}\rangle^{2}_{J,Z_{2}},\end{split}$ (30)
@@
$\begin{split}&\langle C\rangle_{Q,\beta}:=\langle n_{\text{tot}}(n_{\text{tot}}-1)\rangle_{\beta,Q}-\langle n_{\text{tot}}\rangle^{2}_{\beta,Q},\\ &\langle C\rangle_{Z,\beta}:=\langle n_{\text{tot}}(n_{\text{tot}}-1)\rangle_{\beta,Z}-\langle n_{\text{tot}}\rangle^{2}_{\beta,Z},\\ &\langle C\rangle_{Z_{2},\beta}:=\langle n_{\text{tot}}(n_{\text{tot}}-1)\rangle_{\beta,Z_{2}}-\langle n_{\text{tot}}\rangle^{2}_{\beta,Z_{2}}.\\ \end{split}$ (31)
@@ V.2 Contributions from $Q$ and $Z$ for the singularities in the ground-state magic
The results above suggest that in general models, the interplay between the $Q$- and $Z$-parts, or the contributions from the characteristic function and free energy, may result in more intricate behavior of $\tilde{m}_{2}$. The the order of the phase transition plays a significant role here. From this perspective, it is more clear that magic quantified by $\tilde{m}_{2}$ does not necessarily achieve a maximum or minimum at the quantum critical point consequently. This argument is also supported by previous small-scale numerical results for the 2D $\mathbb{Z}_{2}$ gauge theory, obtained using tree tensor network sampling [11]. For the 1D and 2D TFI models studied here, the divergence of the $Q$-part in $[d^{2}\tilde{m}_{2}/dJ^{2}]$, which relates to $\Xi_{P}(\rho)$, suggests a more direct connection between magic and (conformal) criticality. However, this relationship is still subtle, as the $Q$-part may lack a singularity for certain models and general bases.
@@ V.3 Volume-law corrections
To gain deeper insight into the results and many-body magic, we discuss the volume-law corrections of $2$-SRE. In the following discussion, we assume that the ground state exhibits translation symmetry, which holds in many cases including the TFI models considered in this work. If a symmetric ground state exhibits only local magic, for example, |$\ket{\psi}=[(\ket{\uparrow\downarrow}-e^{i\pi/4}\ket{\downarrow\uparrow})/\sqrt{2}]^{\otimes N/2}$, which is a product state of $N/2$ singlet states up to a phase, the strict volume law $\tilde{M}_{2}\propto N$ of the full-state magic follows directly from the additivity property discussed in Sec. II.2. In contrast, if a ground state exhibits nonlocal magic, which cannot be removed via finite-depth local quantum circuits in the thermodynamic limit [87, 88], then nonzero volume-law corrections must appear in the full-state magic. These corrections have similar role with the mutual magic $\mathcal{L}_{AB}(\rho):=\mathcal{M}(\rho_{AB})-\mathcal{M}(\rho_{A})-\mathcal{M}(\rho_{B})$ [9, 61, 15, 62, 63] in a bipartite system ${A}\cup{B}$, where $\mathcal{M}$ is some mixed-state magic monotone. Unlike mutual magic, these volume-law corrections go beyond the bipartition scenario. Although nonzero volume-law corrections are a necessary but not sufficient condition for the presence of nonlocal magic, they serve as a valuable diagnostic tool, particularly given the difficulty of directly identifying nonlocal magic through circuit-based definitions. This mirrors the difficulty of identifying long-range entanglement in many-body systems. For instance, in conformal critical systems, subleading corrections to the area law of entanglement entropy often serve as important indicators [89, 90]. Similarly, we claim that, in the thermodynamic limit, persistent volume-law corrections to SRE at a conformal critical point may indicate nonlocal magic.
To fit the volume-law corrections, we adopt the fitting ansatz
@@
for ground states on a $d$-dimensional lattice with length $L$ with the data of $\tilde{m}_{2}$ in Fig. 2. Additional subleading terms may be present, particularly in higher dimensions. As shown in Fig.  2, the volume-law coefficients $a_{1}$ and $a_{2}$ for the TFI models are close to their corresponding $2$-SRE density, confirming that volume law is exactly the leading term in $2$-SRE.
Fig. 6(a) presents the correction $b_{1}$ for the ground states of the 1D TFI model. Since both the paramagnetic phase and the ferromagnet phase are gapped, in the thermodynamic limit, no nonlocal magic should be captured by $2$-SRE. This reasoning also extends to the 2D case shown in Fig. 6(b). Therefore, we attribute the nonzero values of $b_{d}$ ($d=1,2$) near $J_{c}$ to finite-size effects. More interestingly, these finite-size effects lead to the indication of discontinuities at the critical points. The difference between the two phases are also reflected by the discontinuities of the magical structures reflected by $b_{d}$. This phenomenon contrasts strongly with the full-state magic, from which extracting information of phases and criticalities proves challenging, as we discussed in Sec. V.2. This suggest that the volume-law corrections of magic could provide a stronger diagnostic for criticalities compared to full-state magic. This is especially significant for $\alpha$-SRE, which, as a non-mixed-state measure, fails to provide a reliable or monotonic mutual magic for bipartite systems.
Figure 7: The volume-law scaling $\tilde{M}_{2}=a_{1}L+b_{1}$ fitted for the 2-SRE of the 1D TFI model (PBC) with length $L$ at its critical point $J_{c}=1$. The correction term is determined to be $b_{1}=-0.34(2)$, which is consistent with the theoretical prediction $-\log\sqrt{2}\approx-0.3466$ in Ref. [91].
Importantly, a recent theoretical study has shown that for a (1+1)D conformal critical point, the volume-law correction to the $2$-SRE is a constant [91]. This constant is connected to the non-integer ground-state degeneracy, known as the $g$-factor, of some boundary conformal field theory. In the 1D TFI model, this yields $b_{1,\mathrm{exact}}=-\log\sqrt{2}\approx-0.3466$. To verify this, we carried out low-temperature simulations with $\beta=4L$ and successfully extracted $b_{1}=-0.34(2)$, in excellent agreement with the theoretical expectation, as shown in Fig. 7.
@@ V.4 $2$-SRE of 2D finite-temperature Gibbs states
At the infinite-temperature limit ($\beta=0$), $\tilde{m}_{2}$ is approximately zero within the error bar. For finite temperatures, however, $\tilde{m}_{2}$ exhibits complex behavior, offering minimal valuable information, as shown in Fig. 8(a). Additionally, for the first-order derivative $d\tilde{m}_{2}/d\beta$, we observe a singular point $\beta^{*}\approx 0.687<\beta_{c}$ in the paramagnetic phase, as shown in Fig. 8(b).
@@
In this work, we introduce a novel and efficient QMC method for computing the $\alpha$-SRE and its derivatives in large-scale and high-dimensional quantum many-body systems. Our method is broadly applicable to any sign-problem-free Hamiltonian regardless of the dimension of the system. This versatility makes our method well suited for studying quantum magic across a wide range of physical systems, including those exhibiting spontaneous symmetry breaking, topological phase transitions, topological order, and conformal criticality. As many-body magic remains an underexplored frontier, our algorithm offers a vital computational tool for advancing this area, particularly in light of recent theoretical developments in conformal field theory [91].
@@
We demonstrate our approach on both 1D and 2D ground states, as well as finite-temperature Gibbs states of the TFI model. A key benefit of our approach is that it allows us to isolate the free energy contribution in $\alpha$-SRE and the derivatives, providing deeper understanding into magic and criticality in quantum many-body systems, which are inaccessible before. Particularly, we uncover singularities in the derivatives of the characteristic function contribution in SRE at the conformal critical points in both 1D and 2D. Further, we reveal that the behavior of the $\alpha$-SRE is governed by the interplay between these two contributions, clarifying why full-state magic generally fails to reveal meaningful information about phase structure and criticality, as observed in earlier studies. Meanwhile, the order of the phase transition plays a crucial role for the behaviors of SRE and the singularities in its derivatives. This is quite different from quantum entanglement: for both the 1D and 2D TFI model, previous results show that the quantum entanglement reaches its maximum at the quantum critical points [93, 94].
@@
To further investigate the use of full-state magic, we consider the volume-law corrections, whose nonzero values serve as a necessary condition for the presence of nonlocal magic inherent in system correlations. These corrections, observed in our simulations on finite system sizes, show the evidence of discontinuity at quantum critical points, indicating distinct magical structures on the two sides of the critical point. We consider volume-law corrections to be more significant, and argue that they could be important diagnostics for nonlocal magic at conformal critical points. We leave a more comprehensive exploration of these corrections across a broader class of systems to future work. Finally, we confirm that $\alpha$-SRE is not an effective measure for mixed-state magic in many-body systems by studying an example of finite-temperature phase transition.
In addition, our algorithm shows not only certain magic states of a high-dimensional many-body Hamiltonian can be simulated efficiently on a classical computer, but their magic can be computed efficiently. This underscores the fact that magic is just a necessary but not sufficient resource for quantum advantage.
@@
In closing, we discuss extending our algorithm to compute the magic of a reduced density matrix. The mutual magic $\mathcal{L}_{AB}$ quantifies those magic generates from nonlocal non-Clifford operations, and it could also encode crucial information of many-body systems. Previous studies have shown that the mutual magic $\mathcal{L}_{AB}$ would exhibit stronger signatures than the full-state magic in detecting criticalities and is capable of characterizing certain states in topological quantum field theory [9, 61, 15, 11, 95, 62, 63]. While $\mathcal{L}_{AB}$ is ideally defined using a mixed-state magic monotone, using $\alpha$-SRE to define it, called the mutual stabilizer Rényi entropy (mSRE), has also proven valuable [11, 63, 95], much like quantum mutual information captures entanglement. Its significance has also been theoretically substantiated in the recent development on conformal field theory [91]. Our algorithm can be easily extended to compute reduced density matrix magic by tracing out one subsystem in QMC simulations using the technique in [96, 97, 98, 99], so that the mSRE can be computed. This is also a promising way to study the finite-temperature phase transition and open quantum systems rather than considering the full-state $\alpha$-SRE, as mentioned at the beginning. Our work will enhance the further fusion and intersection of quantum information and many-body physics.
@@ Acknowledgements
We thank P.S. Tarabunga, M. Dalmonte, E. Tirrito, H. Timsina, T. Byrnes, C. Radhakrishnan, K. Warmuz, Z. Liu, G.-Y. Zhu, Y. Tang, and Z.-Y. Wang for helpful discussions. Z.W. acknowledges the China Postdoctoral Science Foundation under Grant No. 2024M752898. The work is supported by the Scientific Research Project (No.WU2024B027) and the Start-up Funding of Westlake University. The authors also acknowledge the HPC centre of Westlake University and Beijng PARATERA Tech Co.,Ltd. for providing HPC resources.start-up funding of Westlake University.